Y=c1e^x+c2e^2x y"-2y+3y=0

1
Expert's answer
2022-02-14T16:45:43-0500

Solution:

To check whether "y=c_1e^x+c_2e^{2x}" is the solution of "y''-2y+3y=0"...(i)

Now, "y=c_1e^x+c_2e^{2x}"

"y'=c_1e^x+2c_2e^{2x}
\\ y''=c_1e^x+4c_2e^{2x}"

Put these in LHS of (i),

"=y''-2y+3y
\\=c_1e^x+4c_2e^{2x}-2(c_1e^x+2c_2e^{2x})+3(c_1e^x+c_2e^{2x})
\\=2c_1e^x+3c_2e^{2x}
\\\ne0"

So, it is not the solution of given D.E.